OpenStudy (anonymous):

Find EA if E(0, -1) and A(1, 2). 2 1.41 3.16 1 i got 1 am i right?

OpenStudy (anonymous):

3.16

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

\[\sqrt{10}\]

OpenStudy (anonymous):

use pythagoras theorem

OpenStudy (anonymous):

vector EA=(1-0)i +(2-(-1))j if i and j are unit vector in the direction of x and y axis. magnitude of EA is square root of ((1^2)+(3^2))=square root of 10=3.16

OpenStudy (anonymous):

thanks you (: